OpenStudy (anonymous):

For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. (hint: you need to find the “k” value)

OpenStudy (anonymous):

The table: X | Y 4 | 10 10 | 25 16 | 40

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

hi

OpenStudy (anonymous):

Hey, @DanJS can you help me out?

OpenStudy (danjs):

varies directly with x, means it is a linear equation in the form y=k * x

OpenStudy (danjs):

put in (x,y) = one of the points and figure out k

OpenStudy (anonymous):

I don't understand where K comes from..

OpenStudy (danjs):

k is the slope, the direct variation of y as x increases

OpenStudy (danjs):

y = (5/4)x

OpenStudy (yamyam70):

\[y = kx , k = y/x\]

OpenStudy (danjs):

5/2x i mean

OpenStudy (danjs):

y = (5/2)x

OpenStudy (anonymous):

@DanJS Isn't there more to the equation that relates to the direct variation?

OpenStudy (danjs):

not here

OpenStudy (danjs):

it is a linear relationship

OpenStudy (yamyam70):

@Answers101 , solving for K From the main formula of direct variation y= kx , solving for K K = y /x When you solve it and you get same results you can conclude that all of them varies directly. Note that direct variation means , if x will rise y will also rise and vice versa, in this case both values of k neither changes.

OpenStudy (anonymous):

So, the answer would be Y = 5/2x that's it?

OpenStudy (anonymous):

@yamyam70

OpenStudy (yamyam70):

Yes you are correct, that equation corresponds to the table given :)

OpenStudy (anonymous):

How do you know that the equation would be 5/2? Because the table shows X shows a decrease of 6 & Y shows a decrease of 15 so would you just simplify 15/6 to get the 5/2?

OpenStudy (yamyam70):

5/2 is not an equation 5/2 is the value of K in the equation y = kx

OpenStudy (danjs):

can you close this out please

OpenStudy (anonymous):

Thanks, @yamyam70

OpenStudy (yamyam70):

@Answers101 alright goodluck :)